Express 2x/ x2-1 in partial fractions and hence find the general solution of the differential equation dy/dx = 2xy / x2 - 1 expressing y explicitly in terms of x.



1
Expert's answer
2021-09-16T00:49:38-0400

The given function "f(x)=\frac{2x}{x^2-1}"

"\frac{dy}{dx}=\frac{2xy}{x^2-1}"

We can write it as,

"\frac{dy}{y}=\frac{2xdx}{x^2-1}"

let "x^2-1= t"

"2xdx=dt"

Hence,

"\frac{dy}{y}=\frac{dt}{t}"

Now, taking the integration of the above equation,

"\int \frac{dy}{y}=\int \frac{dt}{t}"

"\Rightarrow \ln y=\ln t +c"

Now, substituting the value of t

"\ln y = \ln (x^2-1)+c"

Or, we can write it as,

"\Rightarrow \ln y -\ln(x^2-1)=c"


"\Rightarrow \ln\frac{y}{x^2-1}=c"


"\Rightarrow \frac{y}{x^2-1}=e^c"